OpenStudy (chrisneedshelp):

Which of the following is the equation of a direct variation that has a constant of variation equal to -1/2? y=x-1/2 -1/2y=x y= -2x y= -1/2x

OpenStudy (anonymous):

Why it is not B ???

OpenStudy (chrisneedshelp):

ty :P

OpenStudy (anonymous):

@ChrisNeedsHelp can you draw all the answers given???

OpenStudy (unklerhaukus):

@waterineyes why not C

OpenStudy (anonymous):

@ Uncle Rocks let him draw first then I will explain it to you..

OpenStudy (anonymous):

why not D? o.O

OpenStudy (anonymous):

you replied me why not C the same I replied to monokeous..

OpenStudy (chrisneedshelp):

i cannot

OpenStudy (anonymous):

Then tell me is x in the denominator in D option??

OpenStudy (chrisneedshelp):

no

OpenStudy (anonymous):

In B option is y in the denominator??

OpenStudy (chrisneedshelp):

no

OpenStudy (anonymous):

Oh. I see now.

OpenStudy (anonymous):

Ok. :)

OpenStudy (anonymous):

The answer will be D..

OpenStudy (anonymous):

my reasoning is that i followed the y=mx+b

OpenStudy (chrisneedshelp):

thanks guysys

OpenStudy (anonymous):

No that is not the real reason @monokerous

OpenStudy (anonymous):

oh?

OpenStudy (anonymous):

Direct variation goes like this: \[x = k \times y\] and not: \[k \times x = y\] So, According to D; \(y = k \times x\) is right.. But in B: \(k \times y = x\) which is incorrect..

OpenStudy (anonymous):

I see. Thanks for that. I understand now.

OpenStudy (anonymous):

Welcome dear..